Circular No. 6293

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744−28

M. C. Miller, J. M. Quashnock, D. Cole, D. Vanden Berk, D. Q. Lamb, R. C. Nichol and D. Long, University of Chicago, report: “During Jan. 21.539–21.572 UT we carried out optical/near-infrared observations of a region of the sky that included the entire XTE error box (IAUC 6291) for J1744−28, using the Astrophysical Research Consortium’s 3.5-m telescope at Apache Point Observatory. We are confident that we have seen no new object and no significant brightening of any known object in the XTE error box down to a magnitude limit of ∼ 15.5 in g and 17 in r (compared to POSS g and r plates). Similarly, we are confident that we have seen no new object and no significant brightening of any known object in the XTE error box down to a magnitude of ∼ 14 in K′, this compared to a K-band map of the region kindly provided by M. Merrill and I. Gatley (NOAO) and a K′-band map of the field kindly provided by R. M. Catchpole (Glass, Catchpole and Whitelock 1987, MNRAS 227, 373). We are continuing to observe the J1744-28 region each night.”

COMET C/1995 Y1 (HYAKUTAKE) Ephemeris continuation (cf. IAUC 6281) from the orbital elements on

MPEC 1996-B06:

1996 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 23 16^{h}16.^{m}83 −15^{o}42.^{′}5 1.374 1.183 57.^{o}4 44.^{o}5 9.4
28 16 40.99 −12 51.3 1.325 1.148 57.4 46.3 9.2
Feb. 2 17 05.81 − 9 37.8 1.285 1.118 57.2 47.8 9.0
7 17 31.10 − 6 05.6 1.257 1.093 56.8 49.0 8.9
12 17 56.65 − 2 20.0 1.241 1.074 56.3 49.8 8.8
17 18 22.21 + 1 31.9 1.237 1.062 55.6 50.2 8.7
22 18 47.53 + 5 22.6 1.246 1.055 54.9 50.1 8.7
27 19 12.37 + 9 05.0 1.267 1.055 54.1 49.5 8.7
Mar. 3 19 36.51 +12 33.5 1.297 1.062 53.3 48.4 8.8
8 19 59.77 +15 44.2 1.337 1.076 52.5 47.1 8.9
13 20 22.02 +18 35.5 1.384 1.095 51.7 45.4 9.1
18 20 43.16 +21 07.2 1.436 1.120 51.0 43.7 9.3

Total visual magnitudes: Jan. 1.84 UT, 11.3 (Y. Kushida, Yatsugatake South Base Observatory, 0.20-m reflector); 3.54, 10.2 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 14.52, 9.4 (C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector); 17.15, 8.8 (A. R. Baransky, Kiev, Ukraine, 0.11-m reflector); 19.21, 8.9 (W. Hasubick, Buchloe, Germany, 25 × 100 binoculars).

1996 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden